A constant net force of 200 N is exerted to accelerate a car from rest to a velocity of 40 m/s. What is the mass of the cart?


1
Expert's answer
2021-11-16T09:52:25-0500

From the second Newton's law, the mass of the car is:


"m = \dfrac{F}{a}"

where "F=200N" is the net force, "a" is the acceleration of the car. By definition, the acceleration is given as follows:


"a = \dfrac{v-0m/s}{t} = \dfrac{v}{t}"

where "v = 40m\/s" is the final speed of the car, "0m\/s" is its initial speed (start from rest), and "t" is the time required to reach the speed "v".

Thus, obtain:


"m = \dfrac{Ft}{v}"

Answer. "m = \dfrac{Ft}{v}".